OpenStudy (anonymous):

Can someone help me with this: Find all solutions in the interval [0, 2pi) . 7 tan^3x - 21 tan x = 0

OpenStudy (danjs):

try factoring tan(x) out first, so tan(x) = 0, recall identities to replace tan^2 and tan by sin and cos maybe

OpenStudy (danjs):

those probs just play around till you get something in terms of sin and cos using identities and algebra

OpenStudy (danjs):

\[7\tan(x)[\tan^2(x) -3]=0\] \[7\tan(x)[\frac{ \sin^2(x) }{ \cos^2(x) } -3]=0\]

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

On the very left its pie.k and not pie

OpenStudy (anonymous):

@DanJS what are you doing ? :s

OpenStudy (anonymous):

I mean very right* on the bottom +pie.k

OpenStudy (danjs):

lol idk, you have it right, use the unit circle, tan(x)=y/x, remember symmetries if needed for multiple answers over the domain

OpenStudy (anonymous):

Thats better

OpenStudy (danjs):

looks right

OpenStudy (anonymous):

Oh and since its limited between ]0,2pie[ Then you write my answer and add for k=0,1,2